OpenStudy (anonymous):

The midpoint of A (-4, 2) and B(8, 5) is a.(2, 7/2) b.(-6, 3/2) c.(4, 7/2)

OpenStudy (anonymous):

(x1+x2/2, y1+y2/2)

OpenStudy (tyteen4a03):

Remember our mid-point formula: \(x = \frac{x_2 + x_1}{2}\), \(y = \frac{y_2 + y_1}{2}\)

OpenStudy (anonymous):

let us take (x1,y1)=(-4,2) (x2,y2)=(8,5) formula:x1+x2/2,x2+y2/2 -4+8/2=2 y1+y2/2=2+5/2+7/2 so the answer is (2,7/2).

OpenStudy (anonymous):

@turtlemaroon is for sure correct!

OpenStudy (anonymous):

thanks for the medals; glad I could be of help :)